OpenStudy (anonymous):

I need someone who is really good at algebra to help me. I will fan and medal!

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

@WeepingAngel123 PLEASE HELP ME! I will fail if I don't have this and I literally do not understand.

OpenStudy (anonymous):

first know which line equation corresponds to busy street and country road

OpenStudy (anonymous):

Busy street: x-4 country road: -2x+3

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

How you want to find a new line that is similar to busy street but also passes through park

OpenStudy (anonymous):

Now*

OpenStudy (anonymous):

So what part of the line equation for busy street must change and which should remain the same in order to be parallel with park?

OpenStudy (anonymous):

I literally have no clue.

OpenStudy (anonymous):

So you should know first what the meaning behind the equation y=mx+b is. First tell me what does each letter represent on the graph?

OpenStudy (anonymous):

this is slope-intercept form if you want to refer to your class notes

OpenStudy (anonymous):

I should also probably rewrite busy street and country road like this for you: Busy street: y=x-4 Country road: y=-2x+3

OpenStudy (anonymous):

could you just work the first one out for me? I will do the rest.

OpenStudy (anonymous):

Alright. you want a line equation that is parallel to busy road and passes through park. So this means you need a line that has the same slope as busy road but a different y-intercept (same m, different b). So your new equation should look like this: \[y=(1)x+b\] plug in the x and y coordinates that you find from the graph where park is located and solve for b \[4=-8+b\] b=12 is your new y-intercept so the line equation that passes through park and is parallel to busy road becomes \[y=x+12\]

OpenStudy (texaschic101):

Busy street is y = x - 4 the park is at points (-8,4)..x = -8 and y = 4 a parallel line will have the same slope...the slope in y = mx + b form is in the m position. So the slope of y = x - 4 is 1. now we take y = mx + b and sub 1 in for m, the slope. (-8,4)..x = -8 and y = 4 now sub 4 = 1(-8) + b 4 = -8 + b 4 + 8 = b 12 = b so your parallel line is : y = x + 12

OpenStudy (texaschic101):

country road is y = -2x + 3 the house is at points (6,8) a perpendicular line will have the negative reciprocal of the slope given. So our slope is -2. To find the negative reciprocal, " flip " the slope and change the sign. Making the negative reciprocal of -2 = 1/2 now use y = mx + b slope(m) = 1/2 (6,8)..x = 6 and y = 8 now sub 8 = 1/2(6) + b 8 = 3 + b 8 - 3 = b 5 = b so our perpendicular equation is : y = 1/2x + 5

OpenStudy (anonymous):

@texaschic101 you are the best!